I'm making up questions for extra credit in my class. Can you please tell me if this is an answerable question and what you think the answer is? "If the height of a triangle is tripled, the area of the triangle will become...?"
A. 1 1/2 times larger
B. 3 times larger
C. 6 times larger
D. 9 times larger
E. None of the above

Answer by richwmiller(17219)

It is answerable but you should clarify that the base stays the same or that the base also triples.
area = 1/2 b*h
so the area will be triple (three times larger)
example
A=1/2 *2*6
A=6
triple the height to 18
A=1/2*2*18
A=18
18=3*6 and so is triple the original area of 6
Too many choices in your answers.